OpenStudy (anonymous):

find the product of (7xy+3)(4xy-2)

OpenStudy (anonymous):

\[ (7xy+3)(4xy-2)\] requires 4 multiplications

OpenStudy (anonymous):

\[7xy\times 4xy-2\times 7xy+3\times 4xy+3\times (-2)\] is a start

OpenStudy (anonymous):

(7xy + 3) (4xy - 2) = 7xy * 4xy + (-2) * 7xy + 3 * 4xy + 3 (-2)

OpenStudy (anonymous):

= \[28x ^{2}y ^{2} - 14xy + 21xy -6\]

OpenStudy (anonymous):

= \[28x ^{2}y ^{2} + 7x - 6\] This is the answer